OpenStudy (anonymous):

Simplify this expression: √(30)√(6)

OpenStudy (anonymous):

6sqrt5

OpenStudy (anonymous):

\[\sqrt{30}\sqrt{6}=\sqrt{30*6}\] \[=\sqrt{180}\]\[=\sqrt{5*36}\]\[=\sqrt{5*6^2}\]\[=6\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{30}\sqrt{6} = \sqrt{30*6}\] = \[\sqrt{5*6*6}\] = \[6\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{30}\]*\[\sqrt{6}\]=\[\sqrt{180}\] =\[\sqrt{36*5}\] =\[6\sqrt{5}\]